OpenStudy (anonymous):

( (3/4) - (2/3)^2) / ( (3/2) + (7/9) ) Express as a fraction

OpenStudy (saifoo.khan):

Penpel will give it a try.

OpenStudy (anonymous):

11/82????not so sure..ill check my solution

OpenStudy (saifoo.khan):

\[\frac34 - \frac23^2 \div \frac32 + \frac79\] \[\frac34 - \frac49 \div \frac32 + \frac79\]

OpenStudy (anonymous):

I'm kinda confused on what you said saifoo.khan

OpenStudy (anonymous):

(3/4 - 4/9)/ (3/2+7/9) = (27/36 - 16/36)/ (27/18 + 14/18) = (11/36) / (41/18) = ...??!?!? u sure u need to caculate sth this complicated??

OpenStudy (anonymous):

ill continue your solution wamhu.... 11/36(18/41)=11/2(1/41)=11(1)/(41)2=11/82

OpenStudy (anonymous):

thanks...=)

OpenStudy (anonymous):

well it's been very helpfull